OpenStudy (anonymous):

HELP algebra 1 work!

OpenStudy (anonymous):

Complete the chart and determine the solution of the equation f(x) = g(x)

OpenStudy (anonymous):

f(x) = 2x + 24 g(x) = 5(0.5)x

OpenStudy (anonymous):

but none of the choices work which are: x = -2 x = -1 x = 1 x = 20

OpenStudy (anonymous):

@paki help?

OpenStudy (anonymous):

@midhun.madhu1987 help?

OpenStudy (paki):

@Sheraz12345 please help...

OpenStudy (anonymous):

i think the quistion is wrong but im not sure!

OpenStudy (anonymous):

wait what are the 2 equations..

OpenStudy (anonymous):

theyre f(x) = 2x + 24 g(x) = 5(0.5)x

OpenStudy (acxbox22):

set the 2 equations equal to each other

OpenStudy (the_fizicx99):

Mmm, do you mean \(\ \sf g(x) = 5(0.5)^x \) ? If so, because one is linear and one is exponential they intersect when both have the same value http://prntscr.com/4avcl7 therefore it's a x = -2

OpenStudy (anonymous):

OOOOOOOO

OpenStudy (the_fizicx99):

Hence the g(x) = f(x), think of it as 20 = 20

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

now im lost how is it -2?

OpenStudy (acxbox22):

graph the 2 equations and see the x value of where it intersects

OpenStudy (anonymous):

yep its -2

OpenStudy (anonymous):

also i replaced x with -2 and it was 20=20

OpenStudy (the_fizicx99):

20 = 20 is true, so you have to find what "x" coordinate these function have in common that will equal in the same y coordinate.